OpenStudy (anonymous):

I need help with quadratic equations. :3

OpenStudy (anonymous):

When the quadratic equation y = (x − 3)(2x + 7) is written in standard form, what is the value of the coefficient "b"? 2 −21 −1 1

OpenStudy (anonymous):

ax+by+c=0 =>y=2x^2-6x+7x-21 = 2x^2+x-21 coeff b = 1

OpenStudy (anonymous):

Thank you very much!

OpenStudy (anonymous):

no prob !

OpenStudy (anonymous):

Do you want to help me with some more?

OpenStudy (anonymous):

sure till I log off :P

OpenStudy (anonymous):

Ok this one is vertex What is the vertex for the graph of y = −3x2 + 18x − 13? (−3, −40) (−3, −94) (3, 23) (3, 14)

OpenStudy (anonymous):

Substitue the points one by ont in your equation, if you get a 0---well, thats your vertex !

OpenStudy (anonymous):

*one by one

OpenStudy (anonymous):

Ok thanks again. XD